OpenStudy (anonymous):

dy/dx of 3/sinx + cosx ??

OpenStudy (abb0t):

\[\frac{ 3 }{ \sin(x) }+ \cos(x) = 3( \frac{ 1 }{ \sin(x) })+\cos(x)\] 3 is a constant that can be factored out, also note: \[\frac{ 1 }{ \sin(x) } = (\sin(x))^{-1} \neq \arcsin \neq \sin^{-1}(x)\] you can chose either or, using either quotient rule or product rule.

OpenStudy (anonymous):

Thanks

OpenStudy (abb0t):

No problem. Here to help :)